Header Motor, Inc, paid a $3.04 dividend last year. At a constant growth rate of 4 percent, what is the value of the common stock if the investors require a 12 percent rate of return?.
Answer #1
Solution:
value of the common stock    is   $39.52
Working Notes:
Using Gordon growth model : P0 = D1 / (Ke - g), where D1 = D0(1+g)
ke = cost of Equity = required rate of return = 12%
Po=current share price = Value of the common stock = ??
g= growth rate= 4%
D0= Current Dividend=$3.04 per share
P0 = D0(1+g)/(Ke -g)
P0 = 3.04(1.04)/(0.12-0.04)
P0 = $39.52
value of the common stock    is   $39.52
Please feel free to ask if anything about above solution in comment section of the question.